1. Burnham Company collected rent of $3,800 during Year 1. For income tax reporting, the rent is taxed when collected. For financial reporting, the rent is recognized as income in the period earned. At the end of Year 1, the unearned portion of the rent collected during the year amounted to $440. Burnham had no temporary differences at the beginning of the current year. Assume an income tax rate of 30%. What is the amount of the deferred tax asset that should be recognized at the end of Year 1?

2. Exeter Corp. reports warranty expense by estimating the amount that eventually will be paid to satisfy warranties on its product sales. For tax purposes, the expense is deducted when paid. During its first year of operations, Exeter reports pretax accounting income of $100,000. Its income statement includes a $50,000 warranty expense that is deducted for tax purposes when paid in Year 2 in the amount of $30,000 and Year 3 in the amount of $20,000. Exeter is subject to a tax rate of 40%. Prepare the appropriate journal entry to record the company’s income tax expense for Year 1. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

1. Differences = Interest revenue taxed in income tax in year 1 - interest revenue recognised in books in year 1 = $3,800 - $3,360 = $440

Deferred tax asset = $440 * 30% = $132

Therefore, the amount of the deferred tax asset that should be recognized at the end of Year 1 is $132.

2. Differences = $50,000

Deferred tax asset = $50,000 * 40% = $20,000

Income Tax Payable = $150,000 * 40% = $60,000

Income tax Expense = Income tax Payable - Deferred tax asset = $60,000 - $20,000 = $40,000

Journal Entry

Income Tax Expense Dr $40,000

Deferred tax asset Dr $20,000

To Income tax payable $60,000

(Being company's income tax expense recorded for Year 1)